# Probing non-Markovian qubit noise and modeling Post Markovian Master Equation

### v1

@@ I Introduction
In transmon architectures, slow material noise and parasitic ZZZZ couplings generate device-level memory that impacts idle evolution and concurrent gates. Experimental studies report coherence revivals and extra spectral features diagnostic of non-Markovianity; targeted control (tunable couplers, dynamical decoupling) suppresses ZZZZ crosstalk but does not eliminate all memory [13, 14, 15, 4]. A channel-level summary that (i) detects non-Markovianity directly from tomographic data and (ii) models the same data via a physically interpretable kernel $k(t)$ is therefore practically useful.
@@ III.3 Information‑theoretic probe of crosstalk
An intuitive signature of crosstalk is the emergence of correlations between two qubits that were initially uncorrelated. Let the qubits AA and $B$ be prepared in a product state $\rho_{A}\otimes\rho_{B}$. After a joint noise process, the state evolves to a (generically) correlated density matrix $\rho_{AB}(t)$. The total correlation generated can be quantified by the quantum mutual information
@@ IV.1 Quantum State Tomography
where $m$ labels the measurement basis while $k$ labels the outcome. For example, in the single qubit case, the $m$ labels the pauli operators $\sigma_{x},\sigma_{y},\sigma_{z}$ while the $k$ labels the outcome of getting 0 or 11. We will repeat the measurement circuit with sufficiently large number of shots to ensure that the empirical estimation $\hat{p}_{m,k}$ is good enough to approximate the true probability $p_{m,k}$.
@@
where each row of AA is the vectorized POVM element $|E_{m,k}\rangle\!\rangle$, and the matrix multiplication of AA and $x(t)$ is equivalent to applying $\mathrm{Tr}(\rho E_{m,k})$. We first compute the raw least-squares estimate
@@ V.3 Information‑theoretic characterization of qubit crosstalk
A convenient way to quantify correlations generated between two idle qubits AA and $B$ is the quantum mutual information
@@
For two qubits the range is $1\leq S(A|B)\leq 1$. The lower bound for $S(A|B)_{\rho(t)}$ saturates for the maximally entangled state $|\Phi^{+}\rangle_{AB}$, and the upper bound saturates for the product state $\Pi_{A}\otimes\sigma_{B}$ where $\Pi_{A}=\frac{1}{|A|}I_{A}$ (maximally mixed state for system AA) and $\sigma_{B}$ is any valid quantum state for system $B$.
@@ VI Conclusion
Appendix C $Z$ crosstalk effect
@@
    Pure $Z$ coupling. We start with the minimal coherent model in which the main qubit interacts with a single spectator qubit through a $Z$ Hamiltonian. No decoherence is included at this stage.
@@
    $Z$ coupling with local noise. Next, we incorporate independent dephasing and amplitude–damping channels on both qubits to contrast purely coherent and open-system dynamics.
@@
    Multi-spectator scenario with global dissipation. Finally, we generalise to a realistic device layout where the main qubit experiences $Z$ crosstalk with multiple spectators, and every qubit is simultaneously subject to dephasing and amplitude damping.
@@
C.1 $Z$ crosstalk between two qubits
The simplest model of $Z$ crosstalk considers the closed-system evolution generated by the Hamiltonian
@@ C.1 $Z$ crosstalk between two qubits
C.2 $Z$ crosstalk between two qubits with dissipator
We now supplement the coherent $Z$ interaction with two Markovian noise channels—amplitude damping (rate $\gamma_{\downarrow q}$, governing the $T_{1}$ time) and pure dephasing (rate $\gamma_{\phi q}$, governing the $T_{2}$ time). With the standard dissipator
@@ C.2 $Z$ crosstalk between two qubits with dissipator
We see that the $Z$ crosstalk term induces an additional oscillation at frequency $J$. In the opposite limit $\Gamma_{1}/(2J)\gg 1$, where $\Omega\approx\tfrac{\Gamma_{2}}{2}$, one finds
@@
In this regime, the $Z$ crosstalk contribution is negligible, and only purely damped oscillations are observed.
@@
C.3 $Z$ crosstalk with NN spectator qubits
We now generalize the single-spectator case to NN spectators. Consider the $Z$ crosstalk Hamiltonian:
@@ C.3 $Z$ crosstalk with NN spectator qubits
Let us examine how each superoperator acts on the elements of this set. For the qubit detuning and $Z$ crosstalk terms:
@@
Encoding of the NN-qubit parity Hilbert space. To better inspect the structure of the superoperator matrix $\tilde{M}$, we define the computational basis

### v2

@@ I Introduction
In transmon architectures, slow material noise and parasitic $ZZ$ couplings generate device-level memory that impacts idle evolution and concurrent gates. Experimental studies report coherence revivals and extra spectral features diagnostic of non-Markovianity; targeted control (tunable couplers, dynamical decoupling) suppresses $ZZ$ crosstalk but does not eliminate all memory [13, 14, 15, 4]. A channel-level summary that (i) detects non-Markovianity directly from tomographic data and (ii) models the same data via a physically interpretable kernel $k(t)$ is therefore practically useful.
@@ III.3 Information‑theoretic probe of crosstalk
An intuitive signature of crosstalk is the emergence of correlations between two qubits that were initially uncorrelated. Let the qubits $A$ and $B$ be prepared in a product state $\rho_{A}\otimes\rho_{B}$. After a joint noise process, the state evolves to a (generically) correlated density matrix $\rho_{AB}(t)$. The total correlation generated can be quantified by the quantum mutual information
@@ IV.1 Quantum State Tomography
where $m$ labels the measurement basis while $k$ labels the outcome. For example, in the single qubit case, the $m$ labels the pauli operators $\sigma_{x},\sigma_{y},\sigma_{z}$ while the $k$ labels the outcome of getting 0 or $1$. We will repeat the measurement circuit with sufficiently large number of shots to ensure that the empirical estimation $\hat{p}_{m,k}$ is good enough to approximate the true probability $p_{m,k}$.
@@
where each row of $A$ is the vectorized POVM element $|E_{m,k}\rangle\!\rangle$, and the matrix multiplication of $A$ and $x(t)$ is equivalent to applying $\mathrm{Tr}(\rho E_{m,k})$. We first compute the raw least-squares estimate
@@ V.3 Information‑theoretic characterization of qubit crosstalk
A convenient way to quantify correlations generated between two idle qubits $A$ and $B$ is the quantum mutual information
@@
For two qubits the range is $1\leq S(A|B)\leq 1$. The lower bound for $S(A|B)_{\rho(t)}$ saturates for the maximally entangled state $|\Phi^{+}\rangle_{AB}$, and the upper bound saturates for the product state $\Pi_{A}\otimes\sigma_{B}$ where $\Pi_{A}=\frac{1}{|A|}I_{A}$ (maximally mixed state for system $A$) and $\sigma_{B}$ is any valid quantum state for system $B$.
@@ VI Conclusion
Appendix C ZZ crosstalk effect
@@
    Pure ZZ coupling. We start with the minimal coherent model in which the main qubit interacts with a single spectator qubit through a ZZ Hamiltonian. No decoherence is included at this stage.
@@
    ZZ coupling with local noise. Next, we incorporate independent dephasing and amplitude–damping channels on both qubits to contrast purely coherent and open-system dynamics.
@@
    Multi-spectator scenario with global dissipation. Finally, we generalise to a realistic device layout where the main qubit experiences ZZ crosstalk with multiple spectators, and every qubit is simultaneously subject to dephasing and amplitude damping.
@@
C.1 ZZ crosstalk between two qubits
The simplest model of ZZ crosstalk considers the closed-system evolution generated by the Hamiltonian
@@ C.1 ZZ crosstalk between two qubits
C.2 ZZ crosstalk between two qubits with dissipator
We now supplement the coherent ZZ interaction with two Markovian noise channels—amplitude damping (rate $\gamma_{\downarrow q}$, governing the $T_{1}$ time) and pure dephasing (rate $\gamma_{\phi q}$, governing the $T_{2}$ time). With the standard dissipator
@@ C.2 ZZ crosstalk between two qubits with dissipator
We see that the ZZ crosstalk term induces an additional oscillation at frequency $J$. In the opposite limit $\Gamma_{1}/(2J)\gg 1$, where $\Omega\approx\tfrac{\Gamma_{2}}{2}$, one finds
@@
In this regime, the ZZ crosstalk contribution is negligible, and only purely damped oscillations are observed.
@@
C.3 ZZ crosstalk with $N$ spectator qubits
We now generalize the single-spectator case to $N$ spectators. Consider the ZZ crosstalk Hamiltonian:
@@ C.3 ZZ crosstalk with $N$ spectator qubits
Let us examine how each superoperator acts on the elements of this set. For the qubit detuning and ZZ crosstalk terms:
@@
Encoding of the $N$-qubit parity Hilbert space. To better inspect the structure of the superoperator matrix $\tilde{M}$, we define the computational basis